HUBBLE SPACE TELESCOPE DAILY REPORT #1866 PERIOD COVERED: 0000Z (UTC) 04/28/97 - 0000Z (UTC) 04/29/97 Daily Status Report as of 119/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the ultraluminous IRAS galaxy QDOTULIR20. These data will be searched for multiple cores, which imply a merger has occurred, and for starburst regions. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/MA1/S/C 7105 (STIS MAMA Turn-on) The STIS/MA1 was successfully tuned on. This will be the first high-voltage ramp-up of the far-ultraviolet and near-ultraviolet MAMA detectors. The operations was executed as scheduled, and no anomalies were noted. Dark 1.3 Completed Two Sets of NIC1/2/3 7049 (NICMOS Absolute Photometry Test) The Near Infrared Cameras (Cameras One, Two and Three) were used to make observations of standard star P330E using each NICMOS camera at five different wavelengths spanning the NICMOS wavelength range. Biases and an internal flat calabration image were also taken. The observations were completed as planned, and no problems were reported. 1.4 Completed FOC/96 7591 (1997 Spring Heptathlon) The Faint Object Camera (f/96) was used to make calibration observations as part of the Spring Heptathlon. This test is a continuation of the focal plane monitoring program, (Heptathlon). The goal is to measure relative positions of the SIs to within 50 milli-arcseconds, to verify the focal plane plate scale, and by repetition, to assess stability of the focal plane and FGSs. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of STIS/CCD 7092 (CCD Dark Rate and Read Noise) Observations were taken to measure CCD dark current and gather data for making super darks. The observations were completed as planned, and no anomalies were reported. 1.6 Completed WF/PC-2 6447 (Distant Comet Nucleus Sizes) The WF/PC-2 was used to make observations of Comet Shoemaker-Levy-7 (Not to be confused with Comet S-L-9!) in order to study the size of the comet's nucleus. This comet is part of a 10-year, ground-based program to search for physical and chemical differences between the comet's first entering the inner Solar System and after it has been thermally processed. The observations were completed as planned, and no anomalies were reported. 1.7 Completed NIC1/2 7128 (Early Release Observations of Allen's Source in NGC-2264) The Near Infrared Cameras (Both NIC One and Two) were used to make observations of nebula NGC-2264. Observations aim to reveal "Allen's source", at the apex of the "cone", which is heavily obscured in the visible light. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 7553 (Pure Parallel Archive Survey II) WF/PC-2 parallel observations were obtained in the ultraviolet in search for interesting objects. This program is designed to find extragalactic sources with near ultraviolet excesses (3000 A0) outside the galactic plane and far ultraviolet sources (1600 A0) in the galactic plane. The procedure was executed as scheduled, and no anomalies were reported. 1.9 Completed Two Sets of WF/PC-2 6359 (Core Properties of the Bulges of Spiral Galaxies) The WF/PC-2 was used to make snapshot survey observations of spiral galaxies ESO240G12 and NGC-2082 in order to study their central bulges. The aim is to investigate the core properties, correlations with global galactic properties, and the frequency of occurrence of multiple and/or miscentered nuclei. The observations were completed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 7029 (WF/PC-2 Cycle 6/SMOV Lyman Alpha Check-Part 2) The WF/PC-2 was used to make observations of the WD0310-688 in order to obtain far-ultraviolet (Lyman Alpha) measures of the system's throughput. This is to quantify any contamination of the WF/PC-2 pick-off mirror caused by SMOV activities. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: There were no reported operational problems. The memory monitors for MAMA high voltage turn-on were set-up at 118/1344Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH